OpenStudy (anonymous):

Find derivative of sqrt(x^2 + 2x - 1)

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

Can you show all the steps?

OpenStudy (babyslapmafro):

This might help as well...\[\sqrt{x^2+2x-1}=(x^2+2x-1)^{1/2}\]